### v1

# ARC: Argument Representation and Coverage Analysis for Zero-Shot Long Document Summarization with Instruction Following LLMs

Abstract
Integrating structured information has long improved the quality of abstractive summarization, particularly in retaining salient content. In this work, we focus on a specific form of structure: argument roles, which are crucial for summarizing documents in high-stakes domains such as law. We investigate whether instruction-tuned large language models (LLMs) adequately preserve this information. To this end, we introduce Argument Representation Coverage (ARC), a framework for measuring how well LLM-generated summaries capture salient arguments. Using ARC, we analyze summaries produced by three open-weight LLMs in two domains where argument roles are central: long legal opinions and scientific articles. Our results show that while LLMs cover salient argument roles to some extent, critical information is often omitted in generated summaries, particularly when arguments are sparsely distributed throughout the input. Further, we use ARC to uncover behavioral patterns—specifically, how the positional bias of LLM context windows and role-specific preferences impact the coverage of key arguments in generated summaries, emphasizing the need for more argument-aware summarization strategies.
1 Introduction
LLMs have made remarkable progress in text summarization, often generating summaries preferred by human evaluators in domains like news Zhang et al. (2024); Liu et al. (2024b). However, summarization in structured, highly informative domains presents unique challenges that remain underexplored. One prominent challenge is preserving salient argument roles in the generated summaries, which can be challenging as arguments can be sparsely distributed across the input Elaraby and Litman (2022). The ability of LLMs to selectively retain argument roles is therefore a crucial test of their utility in generating reliable summaries in high-stakes domains.
Prior work in legal summarization has shown that explicitly modeling argument roles—either during finetuning Fabbri et al. (2021a); Elaraby and Litman (2022) or through post hoc re-ranking Elaraby et al. (2023)—improves the coverage of critical argumentative content. These findings suggest that pretrained language models (PLMs) may struggle to capture structured discourse elements such as arguments without targeted supervision. However, it remains an open question whether instruction-tuned LLMs, trained with broad and often general-purpose supervision, can inherently identify and preserve salient argumentative information without explicit signals about saliency or additional tuning. In this work, we go beyond conventional summary evaluation metrics such as fluency, factuality, and coherence to address a core question: Do LLMs effectively prioritize and preserve the most salient argumentative content in their summaries?
To address this question, we introduce Argument Representation Coverage (ARC), a framework for evaluating how well LLM-generated summaries capture salient arguments. ARC measures coverage at three levels of granularity: (1) Argument Set Coverage (), assessing collective coverage of full salient argument set; (2) Independent Role Coverage (), evaluating each argument role separately as atomic units; and (3) Subatomic Coverage (), examining fine-grained factual units within roles.
Figure 1 presents an example of how ARC operates across multiple levels of abstraction. While provides a single holistic score capturing overall argument coverage, it fails to offer fine-grained insight into which specific arguments are preserved or omitted—information critical for downstream analyses such as bias detection. addresses this by assessing the preservation of each argument role independently, although it considers an argument entirely unsupported if even one element is missing. To mitigate this limitation, decomposes arguments into subatomic factual units, enabling a more continuous and informative scoring mechanism. This multi-level approach is designed to overcome the limitations of coarse Likert-scale ratings, which have been shown to suffer from low inter-annotator agreement in prior work Elaraby et al. (2023); Krishna et al. (2023a).
Prior work highlights that LLMs often exhibit positional biases—favoring content from the beginning or end of long documents Liu et al. (2024a); Ravaut et al. (2024a); Wan et al. (2024). Furthermore, it remains unclear whether LLMs also favor certain types of salient information that share the same structure (e.g., argument roles) during summarization. Using ARC, we conduct two key analyses: (1) How does the position of arguments in the source document affect their inclusion in summaries? and (2) Are certain argument roles disproportionately favored over others? The latter analysis is crucial, as uneven role coverage may lead to biased or incomplete summaries that misrepresent the original discourse. We conduct our experiments across two domains where argument structure plays a central role in understanding the essence of the document: Long Legal Opinions and Scientific Articles.
Our contributions are in two folds: (1) we propose ARC, a multi-granularity evaluation framework for evaluating argument coverage; and (2) we present a systematic analysis of how positional and role-specific biases affect LLMs’ ability to preserve salient argumentative content in summaries.
| Dataset | # Docs | Input Length | Summary Length | % Roles in Input | % Roles in Summary |
| CANLII | |||||
| DRI | - |
2 Related Work
Information Saliency in LLMs.
Content selection remains a core challenge in summarization. Trienes et al. (2025) found weak alignment between LLMs’ saliency preferences and human judgments. While LLMs can produce summaries preferred over human references in domains like news Zhang et al. (2024); Liu et al. (2024b), they still benefit from content planning. For example, Adams et al. (2023) showed that planning entity mentions improves the information density in GPT-4 generated summaries at the same summary length when compared to summaries generated without entitiy planning. We extend this line of work by treating argument roles as a structured form of saliency and analyzing their preservation in LLM-generated summaries.
Limitations of LLMs in Long-Document Summarization.
LLMs face persistent issues when summarizing long texts, notably the U-shaped positional bias—favoring content at the beginning and end while neglecting the middle Ravaut et al. (2024b). This leads to degraded faithfulness in long-form outputs Wan et al. (2024). We expand this analysis by quantifying how positional bias affects the coverage of salient argumentative content.
Argument Mining and Abstractive Summarization.
Incorporating argument structures into summarization has shown promise across domains, including dialogues Fabbri et al. (2021a), legal texts Xu et al. (2020, 2021); Elaraby and Litman (2022); Elaraby et al. (2023) and scientific documents Fisas et al. (2016). We build on this by assessing whether instruction-tuned LLMs can cover salient arguments without the external argument role information, particularly in summarizing legal and scientific texts.
Evaluation Metrics for Long-Form Summarization.
Standard metrics often fall short in reflecting human preferences, especially for long documents Fabbri et al. (2021b); Krishna et al. (2023b). To improve reliability, recent work has introduced unit-based metrics—such as Atomic Content Units (ACUs) Krishna et al. (2023a) and structured factuality scores Min et al. (2023); Yang et al. (2024)—to reduce subjectivity. Extending this idea, we propose ARC, which uses argumentative structures as evaluation units and introduces subatomic granularity to assess fine-grained argument coverage.
3 Datasets111More analysis and examples in Appendix A.
We employ two datasets that include both argument role annotations and reference summaries: CANLII Xu et al. (2021), representing the legal domain, and DR. INVENTOR (DRI) Fisas et al. (2016), representing the scientific domain. An overview of dataset statistics is presented in Table 1. Both datasets consist of long-form documents paired with long-form reference summaries that average words Krishna et al. (2023b).
| Dataset | Argument Role | Example |
| CANLII | Issue | Damage to both vehicles exceeded the insurance deductibles and both parties claim damages against each other. |
| Conclusion | Fault for this accident was attributed to the defendant and to the plaintiff. | |
| Reason | Jurisdictional error is not to be equated with error of law. | |
| DRI | Own Claim | Semi-Lagrangian contouring offers an elegant and effective means for surface tracking with advantages over competing methods. |
| Background Claim | Accurate modeling of human motion remains a challenging task. | |
| Data | Animation is constrained due to hardware constraints. |
3.1 Legal Opinions: CANLII
The CANLII dataset consists of legal cases annotated at the sentence level for argument roles. Notably, only of the input text is labeled with argument roles, yet these argumentative sentences account for of the reference summaries (Table 1). This substantial mismatch highlights a haystack-like challenge: models must accurately identify and prioritize the sparse yet highly salient argumentative content when generating summaries.
Argument roles in CANLII are annotated using the IRC scheme Xu et al. (2021), which categorizes roles into three types: Issues: Legal questions raised in the case. Reasons: Justifications provided for judicial decisions. Conclusions: Final rulings addressing the identified issues. These annotations enable a fine-grained evaluation of argument coverage in generated summaries. Examples of IRCs are shown in Table 2.
3.2 Scientific Articles: DRI
The DRI dataset consists of computer graphics articles, each annotated at the sentence level for rhetorical roles and paired with three human-written summaries. Notably, these rhetorical roles are not necessarily argumentative, making it challenging to assess argument coverage. To address this limitation, the extended version of the dataset, SCI-ARG Lauscher et al. (2018), enriches the DRI annotations by incorporating argument role annotations and their relations. These annotations follow a modification of the -argument roles described in Toulmin model Toulmin (2003), by reducing them into three types: Own Claim: Sentences that directly support the author’s central argument. Background Claim: Sentences that reference prior research or established domain knowledge. Data: Empirical evidence that supports or refutes claims, such as experimental results or literature citations. An example of each role is shown in Table 2.
Since the argument annotations are span-based, we map them back to complete sentences using lexical matching assigning the sentence with argument role spans if of its words falls within the sentence boundaries. A motivating feature behind selecting this corpus for our analysis is the sentence-level annotation for relevance scores on a Likert scale , which indicates the degree of relevance to the summary. A relevance score of signifies that the sentence is "relevant to the summary," while a indicates it is "very relevant to the summary." In our evaluation, we focus on argument role coverage for sentences with argument roles and a Likert score of (indicating high relevant argument roles to the summary). Table 1 shows that unlike legal opinions, where argument roles are sparsely distributed, scientific articles contain argumentative content throughout the document, posing a challenge of selectivity rather than retrieval. In DRI, sentences that contain at least 1 argument role account for of the input text (shown in Table 1). Although the dataset does not provide gold-standard summaries annotated for argument roles, we analyze the sentences with a Likert score of based on their argument role annotation. Among these sentences, contain at least one argument role, reinforcing the strong connection between argument roles and summarization relevance in this domain.
4 The ARC framework
4.1 Overview and Notations
Given a generated summary and a set of salient arguments , we define a coverage function , where: . Here, represents the evaluation unit against , where: . represents the full set of salient arguments in the document. is an individual argument . is a subatomic factual unit, where and is the set of atomic facts derived from argument .
Full Argument Set Coverage
Following Elaraby et al. (2024), we define based on a Likert-scale annotation (1-4), which is then normalized to a range:
where is the Likert score assigned to the argument set coverage.
Independent Argument Role Coverage
Each argument is evaluated independently as:
Subatomic Argument Coverage
An argument is further decomposed into subatomic fact units , where for :
The scores obtained by at different granular levels provide a deeper assessment of model behavior on the argument level. Table 3 shows ARC scores computation for an number of salient arguments.
| Granularity | Computation |
| Full Set | |
| Individual Roles | |
| Subatomic Units |
| Metric | Metric Type | Expert 1 | Expert 2 | Average | |||
| ROUGE-1 | Lexical | 0.3455 | 0.4809 | 0.3382 | 0.4495 | 0.3757 | 0.5318 |
| ROUGE-2 | 0.3233 | 0.4150 | 0.2860 | 0.4132 | 0.3291 | 0.4734 | |
| ROUGE-L | 0.2824 | 0.3874 | 0.3463 | 0.4455 | 0.3418 | 0.4764 | |
| BERTScore | Semantic | 0.3187 | 0.4092 | 0.2921 | 0.3977 | 0.3344 | 0.4756 |
| (sent) | Entailment | 0.3676 | 0.5157 | 0.3567 | 0.4368 | 0.3204 | 0.4654 |
| (doc) | 0.4512 | 0.4894 | 0.2617 | 0.3438 | 0.3747 | 0.4758 | |
| (sent) | 0.3204 | 0.4654 | 0.3567 | 0.4368 | 0.3676 | 0.5157 | |
| (doc) | 0.4512 | 0.4894 | 0.2617 | 0.3438 | 0.3747 | 0.4758 | |
| LLM-Judge | 0.6713 | 0.7288 | 0.4072 | 0.4793 | 0.5867 | 0.6898 | |
| 0.4884 | 0.6034 | 0.3453 | 0.4527 | 0.4474 | 0.5971 | ||
| 0.5806 | 0.7023 | 0.5135 | 0.6353 | 0.6025 | 0.7560 | ||
| 0.5025 | 0.6026 | 0.4593 | 0.5727 | 0.5142 | 0.6642 | ||
| 0.5213 | 0.6507 | 0.4347 | 0.5304 | 0.5202 | 0.6959 | ||
4.2 Choice of
Given the limited availability of large, diverse argument coverage datasets, we use the annotated dataset from Elaraby et al. (2024), comprising legal opinions from the CANLII dataset with corresponding summaries. Each summary is annotated for argument coverage using a 4-point Likert scale at the full argument set level.333See Appendix B for scale definitions.
Computing using LLM-judge
We use GPT-4o with temperature sampling as an automated evaluator, following prior work that demonstrated its strong correlation with human judgments in summarization tasks Liu et al. (2023b).
: We prompt the model to assign a Likert score (1-4) based on coverage guidelines from Elaraby et al. (2024). Scores are then normalized to for comparability with other metrics.
: The model assigns binary scores (1 = fully supported, 0 = missing or inconsistent) for each argument .
: Building on recent factuality frameworks that decompose summaries into atomic fact units for grounded evaluation Min et al. (2023); Lee et al. (2024); Tang et al. (2024); Yang et al. (2024), ARC instead decomposes key arguments need to be retained. This shift focuses evaluation on the completeness of argument roles rather than factuality evaluation. Unlike prior completeness work, which treats all information uniformly, argument roles impose a structured prioritization, which is sparse in legal cases and dense in scientific articles, enabling more targeted analysis to indicate if it’s still worthy including structured information such as argument roles for a more complete summary.
Following the decomposition method of Yang et al. (2024), each argument is decomposed into a set of factual units using GPT-4o. Units are filtered by a fine-tuned entailment model (DeBERTa; He et al. )444https://huggingface.co/cross-encoder/nli-deberta-v3-base to keep only entailed facts555Decomposition details in Appendix C.. Each unit is then verified against the summary , yielding , where indicates support and specifies the error type (missing or non-factual). This enables fine-grained argument coverage evaluation while distinguishing hallucination from information loss666Evaluation prompts in Appendix D..
Reducing LLM-Judge Cost Evaluating argument coverage at scale using GPT-4o is costly, particularly for the legal dataset (CANLII), where each document contains numerous argument units and their subatomic facts. To mitigate this cost, we train classifiers to approximate GPT-4o judgments, thereby reducing reliance on LLM-based evaluation.
We sample cases from CANLII and use GPT-4o to compute and for all models. This data serves as the training set for two classifiers. The first classifier, , is a binary classifier that predicts argument-level support, trained to approximate the judgments of GPT-4o such that . The second classifier, , is a three-way classifier that predicts whether a subatomic fact is supported, missing, or non-factual, ensuring that .
We explored several models, including DeBERTa (base and large), LegalBERT Chalkidis et al. (2020), and BigBirdRoBERTa Zaheer et al. (2021). Among all models, the DeBERTa-large demonstrated the highest performance, achieving an F1-macro score of for subatomic facts evaluation and for role evaluation against summary predictions, using -fold cross-validation 777Training setup and evaluation against GPT-4o are provided in Appendix E..
ARC Scores Against Human Evaluation We compare ARC scores with lexical and semantic metrics from Elaraby et al. (2024), as well as entailment-based metrics, particularly SummaC Laban et al. (2022) (both zero-shot and convolution-based), which measures alignment between source articles and summaries. For entailment evaluation, argument roles are treated as the hypothesis and generated summaries as the premise, with SummaC computed at both document- and sentence-level granularities. Table 4 shows that ARC consistently outperforms lexical, semantic, and entailment metrics in correlating with expert annotations, with one exception: for against Expert 2. The lower performance of suggests that treating a full argument as an atomic unit introduces excessive penalization for minor omissions and inconsistencies, leading to misalignment with holistic human judgments. In contrast, achieves the highest correlations with Expert 2 and the overall expert average, enabling getting a nuanced and interpretable score without losing the holistic overall coverage.
When using trained classifiers, we observe a drop in performance but an improvement in , which is likely due to distributional shifts between training and human-evaluated summaries. Nonetheless, the stricter nature of appears to amplify small deviations, reducing alignment with expert assessments. Given its strong expert correlation and interpretable error predictions, we adopt ARC as the primary metric for coverage evaluation, providing a reliable and fine-grained lens on model performance.
4.3 Obtaining generated summaries
To handle the long-form nature of argumentative texts, we employ open-weight LLMs with extended context windows: LlaMA-3.1-8B-instruct (LlaMA-3.18-B) Grattafiori et al. (2024), Mistral-8B-instruct(Mistral-8b) Jiang et al. (2024), and Qwen2-7B qwe (2024). For Qwen2-7B, to support context lengths exceeding 32k tokens, we integrate YARN embeddings Peng et al. into the model configuration prior to deployment. Inference is conducted using VLLM Kwon et al. (2023) for scalability. Summaries are generated using temperature sampling, capped at tokens to ensure fair long-form generation. Each document is prompted with: "Read the following text and summarize it: {input document}. Summarize in {reference summary word length} words. Summary:". The target length is dynamically set to match reference summaries for comparability. For DRI, the length is fixed to the longest reference summary to encourage maximal argument retention.
5 Results and Analysis
5.1 Do LLMs cover salient arguments effectively?
We compute ARC888Full ARC and ARC results are in Appendix F. scores for human reference summaries on both the CANLII and DRI benchmarks using GPT-4o. For CANLII, this evaluation serves to assess the robustness of the metric by verifying whether it assigns near-perfect scores to ideal human-written summaries. For DRI, the goal is to measure how much human references cover salient argument roles compared to LLM-generated summaries. As shown in Figure 2, argument coverage remains imperfect across all models, with Mistral-8B lagging behind both LlaMA-3.18B and Qwen2-7B. Coverage for DRI is consistently higher and close to the reference summary compared to CANLII across all evaluated models. This pattern highlights the greater challenge of preserving salient argumentative information in legal texts, where arguments are often sparsely distributed across lengthy and complex contexts.
Error types We analyze the types of error extracted from of the ARC evaluation. Figure 3 shows that the most frequent error is missing, indicating that facts in arguments are often omitted rather than misrepresented. While factual inconsistencies exist, they are less prevalent than missing information. These findings emphasize that beyond hallucination, ensuring comprehensive salient information coverage remains a critical challenge in summarization.
5.2 Do argument positions in the source affect their coverage in summaries?
Positions of arguments in the source from the perspective of LLMs
Following Ravaut et al. (2024b), we start by analyzing the positions where included LLMs look at in its context window. We leverage the lexical greedy approach for source sentences identification Ravaut et al. (2024a); Adams et al. (2023) by iteratively adding sentences in the source that maximizes ROUGE-1 score until there is no further improvement. We analyze the source sentence indices by their argument role annotations.
Figure 4 reveals a distinct U-shaped context window across all models in both datasets, with the effect being particularly pronounced in CANLII. Analyzing source sentences based on their argument role annotations suggests that argument positions are strongly influenced by this U-shaped pattern. This is especially concerning for CANLII, where reference summaries indicate that arguments do not adhere to a fixed positional pattern. In contrast, reference summaries in DRI more closely align with the LLM context window distribution.
Effect of context window positional bias on coverage
We analyze the positions of salient arguments using . For CANLII, we apply greedy sentence selection while restricting the selection to the annotated arguments in both the reference summary and the input document, thereby reducing computational cost and ensuring that only arguments in the reference summary are mapped to arguments in the input. For DRI, we directly select the positions of arguments in the input document with a relevance score . We follow Ravaut et al. (2024b) by computing the mean relevant position of salient arguments and measuring the Pearson correlation between this mean and .
| Model | CANLII | DRI |
| LlaMA-3.1-8B | -0.230 | 0.129 |
| Mistral-8B | -0.369 | -0.055 |
| Qwen2-7B | -0.301 | -0.223 |
Table 5 999Correlation with full ARC scores included in Appendix G. shows a significant negative correlation in CANLII (), indicating that LLM context windows impact argument coverage. In DRI, correlations are weaker and not significant, with LlaMA-3.1-8B even showing a slight positive trend. These patterns align with Figure 4: reference summaries in CANLII differ in distribution from generated ones, while in DRI, argument roles better reflect the U-shaped context window bias. This suggests that context-window positional bias can hinder argument coverage, especially when arguments are sparsely distributed.
5.3 Do LLMs cover argument roles disproportionally?
We propose a bias score . This score is grounded in the metric. We compute for each argument role by first calculating its score and normalizing it by the role’s prior frequency in the source document. This adjustment corrects for overrepresented roles, ensuring the bias score reflects true disparities in coverage. The final bias score for an argument role is defined as:
Here, denotes the frequency of argument role in source document , and is the total number of arguments in . To mitigate bias from argument length and position—especially in CANLII, where longer roles and mid-position arguments negatively affected coverage—we compute within groups allowing up to word-length variation. We also control for position by selecting CANLII articles in which at least of arguments appear within the first or last of the case.
Figure 5 101010Appendix H includes bias analysis without prior frequency normalization to avoid denominator inflation. shows that salient arguments are not covered equitably across roles. In CANLII, conclusions are consistently better covered than issues and reasons across all groups. While length control reduces some role-related bias, conclusion remain the most covered, even under both length- and position-controlled settings. In contrast, for DRI, own claims are less covered relative to other roles, but when compared to reference summaries, their coverage is comparable. This suggests the lower bias score arises from normalization, reflecting the overrepresentation of own claims in the source corpus rather than a true coverage gap. Finally, while there is a bias to human reference in case of CANLII , the gap in argument coverage is notably larger for LLM outputs compared to DRI.
6 Conclusion and Future Work
We introduced ARC, a novel evaluation framework for evaluating how well LLM-generated summaries preserve salient arguments. Our multi-level formulation—spanning full-set, role-level, and subatomic coverage—correlates more strongly with human judgments compared to standard lexical, semantic, and entailment-based metrics, with subatomic evaluation showing highest overall correlation with human judgments. Using our proposed metric, we identified two key limitations in summarizing long legal opinions. First is positional bias, where LLMs tend to overrepresent content at the beginning or end of the input context, affecting the coverage of sparsely distributed arguments. Second is role bias, where models disproportionately favor covering conclusions over other roles such as issues and reasons. Future work can build on this work by exploring incorporating explicit argument representations during training or during prompting LLMs for a more informative summaries in high stakes domains.
Limitations
While the ARC framework enables a comprehensive multi-level evaluation of argument coverage, several limitations remain. First, our evaluation is constrained by the lack of dedicated benchmarks explicitly designed for argument coverage. Existing datasets offer limited annotation granularity, particularly at the subatomic level, which restricts the reliability of fine-grained assessments. As a result, our decomposition into atomic argument units depends on LLM-based prompting and entailment filtering, both of which may introduce inaccuracies. Second, the relatively small size of the DRI dataset (40 documents) limits the generalizability of our findings in the scientific domain. While we were constrained by the available datasets, developing larger, rigorously annotated datasets is an important direction for future work, though it is beyond the scope of this study. Finally, due to computational constraints, our evaluation only includes models that fit within available memory resources. We encountered out-of-memory issues with several larger models. Future work should consider incorporating larger-scale models to determine whether argument coverage improves with increased model capacity or if the observed limitations persist.
Ethics Statement
Our study complies with the ACL Ethics Policy. We primarily evaluate academically available datasets designed explicitly for research purposes,which we obtained through license agreement with the authors of both datasets, thus minimizing privacy risks. Additionally, our work acknowledges potential biases and inaccuracies inherent to LLM-generated outputs, including misrepresentation or omission of critical information from summaries, which could have significant implications in high-stakes domains such as law and science. Researchers and practitioners utilizing our framework should exercise caution and validate results carefully before applying these models in sensitive or consequential decision-making contexts.
Appendix A Extended analysis on included datasets
A.1 Examples from included datasets
Table 6 presents an excerpt from a legal opinion in the CANLII dataset, with arguments highlighted in both the input and the reference summary. Table 7 provides an excerpt from a scientific article in the DRI dataset, with highlighted arguments in the input. Although the documents are truncated for space, the examples clearly illustrate a key distinction: in CANLII, arguments constitute a smaller fraction of the input, whereas in DRI, the input is densely populated with argumentative content.
| Input Article (Truncated) |
| Q.B. A.D. 1987 No. CS 1159 J.C.R., Regina. Applicants seek to quash a search warrant issued by a justice of the peace. The respondent, a justice of the peace, issued a search warrant to search a dwelling house for weapons allegedly used in an attempted armed robbery. The applicants claim the warrant was unlawfully issued without proper grounds. Specifically, the sworn information relied solely on hearsay from an unidentified informant, lacking corroborating details. The applicants argue that no reasonable or probable grounds were disclosed to believe the weapons would be found at the searched location. They highlight that the informant’s reliability was not established, nor was there an oath affirming the informant’s credibility. The search warrant was issued under Section 443(1)(b) of the Criminal Code, which allows a justice to issue a warrant if reasonable grounds exist to believe evidence of an offence will be found. The court explains that on applications to quash a warrant, the reviewing judge cannot substitute their opinion for that of the justice of the peace. Instead, the judge must simply determine whether any evidence existed upon which the justice could be satisfied on reasonable grounds. Reliance on confidential informants is permitted, even if detailed particulars are absent, provided sufficient basis exists for reliability. Past cases (e.g., Re Lubell, Re Dodge) have accepted similar levels of disclosure to protect informant anonymity. The court notes that substantial compliance with Section 443 is sufficient; perfection in drafting is not required. Given practical constraints faced by peace officers preparing information, reasonable latitude must be given in interpreting the sworn information. The judge concludes that although more information could have been provided, there was sufficient evidence upon which the justice could reasonably issue the warrant. Accordingly, the respondent acted within her jurisdiction, and the application to quash the warrant is dismissed. |
| Reference Summary |
| Warrant issued to search a dwelling house for weapons allegedly used in an attempted armed robbery. The affidavit in support referred to an unknown informant. Judge applied the test that the justice of the peace ‘must be satisfied on reasonable grounds.’ Substantial compliance found and warrant upheld. |
| Input Article (Truncated) |
| Our method maintains an explicit polygonal mesh that defines the surface, and an octree data structure that provides both a spatial index for the mesh and a means for efficiently approximating the signed distance to the surface. At each timestep, a new surface is constructed by extracting the zero set of an advected signed-distance function. Semi-Lagrangian backward path tracing is used to advect the signed-distance function. One of the primary advantages of this formulation is that it enables tracking of surface characteristics, such as color or texture coordinates, at negligible additional cost. We include several examples demonstrating that the method can be effectively used as part of a fluid simulation to animate complex and interesting fluid behaviors. The fundamental problem of tracking a surface as it is advected by some velocity field arises frequently in applications such as surface reconstruction, image segmentation, and fluid simulation. Unfortunately, the naive approach of simply advecting the vertices of a polygonal mesh quickly encounters problems such as tangling and self-intersection. Instead, a family of methods, known as level-set methods, has been developed for surface tracking. These methods represent the surface implicitly as the zero set of a scalar field defined over the domain. Level-set methods avoid dealing with topological changes but require high-order conservation law solvers. In contrast, our method constructs a surface directly using semi-Lagrangian contouring without solving PDEs, preserving surface detail efficiently. Using adaptive octree data structures, we can efficiently and reliably construct the new surface and corresponding signed-distance function. This allows tracking surface properties such as color or texture coordinates directly on the polygonal mesh during advection, enabling realistic animation of complex fluids. Prior methods often suffered from volume loss and smoothing artifacts, particularly in underresolved, high-curvature regions. By using an explicit surface representation, we compute exact distances near the mesh and avoid substantial interpolation errors. … Finally, the method produces detailed, flicker-free animations of fluid behavior, demonstrating significant advantages over traditional level-set and particle-based approaches. |
| Reference Summary |
| This article presents a semi-Lagrangian surface tracking method that explicitly represents the surface as a set of polygons. The new surface and corresponding signed-distance function can be efficiently and reliably constructed using adaptive octree data structures. One of the primary advantages of this method is that it enables tracking surface characteristics, such as color or texture coordinates, or even simulation variables, accurately at negligible additional cost. These properties can be easily stored directly on the polygonal mesh and efficiently mapped onto the new surface during semi-Lagrangian advection. At each timestep, a new surface is constructed by extracting the zero set of an advected signed-distance function. The explicit representation provides advantages on computing exact signed-distance values near the mesh and storing properties on mesh vertices. It also facilitates other common operations developed for manipulating and rendering triangle meshes. To avoid the topological difficulties of directly updating an explicit surface representation, the surface is updated in time through an implicit representation. The implicit representation is then used to construct a new mesh and extracted using a contouring algorithm. For its simplicity, robustness, and speed, marching-cubes method is used for contouring. After the triangle mesh has been extracted, true distance values are assigned to the vertices of octree. This process is known as redistancing, which comprises three steps: coarsen the octree; compute exact distances at vertices; run a fast marching method over the remaining vertices. Finally, this method is able to produce detailed, flicker-free animations of complex fluid motions. |
A.2 Distribution of arguments across the input
Figure 6 illustrates the distribution of argument roles across the source documents. In CANLII, Conclusion statements predominantly appear toward the end of the document, while Issue statements are concentrated near the beginning. In DRI, Background claims are more frequent at the start of the document, which aligns with the conventional structure of scientific writing where literature reviews—typically containing claims from prior work—are introduced early on.
A.3 Distribution of salient arguments
Figure 7 presents the distribution of argument roles in CANLII reference summaries and in DRI sentences annotated with a relevance Likert score of 5 (indicating very high likelihood of inclusion in a summary). In CANLII, the distribution of argument roles is relatively balanced across categories, whereas in DRI, own claims—statements made directly by the authors—dominate the content.
A.4 Rhetorical roles per relevance to summary likert score
To better understand the relationship between rhetorical structure and relevance to the summary, we compute the percentage of each rhetorical role across Likert-rated sentences in the DRI corpus. As shown in Figure 8, non-argumentative content dominates among sentences rated as totally irrelevant to the summary (Likert score 1). However, as the perceived relevance increases, argumentative content becomes more prominent, with Own Claim consistently emerging as the most frequent rhetorical role across all higher-quality categories. This trend highlights a clear shift toward structured argumentative writing in more relevant reflections.
Appendix B Likert scores based on human evaluation
Table 8 shows the Likert scale from to definitions.
| Rating | Explanation of the Generated Summary Rating Scale |
| 1 | No arguments covered: The generated summary did not cover the highlighted arguments in the reference summary or covered them only inadequately. |
| 2 | Few arguments covered: The generated summary adequately covered only a limited number of the highlighted arguments in the reference summary. |
| 3 | Most arguments covered: The generated summary adequately covered most of the arguments highlighted in the reference summary. |
| 4 | All arguments covered: The generated summary adequately covered all the highlighted arguments in the reference summary. |
Appendix C Fact decomposition algorithm
Algorithm 1 outlines the decomposition process for an arbitrary argument , performed via prompting GPT-4o.
| Prompt Given to GPT4-o for Argument Decomposition |
| Task: |
| Extract a set of atomic facts—statements that can be directly inferred from the argument without interpretation, assumptions, or redundancy. |
| Guidelines: |
| • Extract only explicitly stated atomic facts. • Do not repeat facts or infer from external knowledge. • Maintain granularity: each fact should be minimal yet complete. • Output a valid Dictionary object where each key is "fact1", "fact2", etc., and the values are the corresponding atomic facts. • No additional text or formatting; dictionary object only. • Each argument must yield at least one atomic fact. |
| Example Output Format: |
| "fact1": "First atomic fact", |
| "fact2": "Second atomic fact", |
| "fact3": "Third atomic fact" |
| Input: |
| {argument} |
| Output: |
| (Dictionary object only) |
| Argument (Issue): |
| FIAT. The father applied to have the mother cited for contempt for denial of access. |
| ✓ Fact 1: The father applied to have the mother cited for contempt. (Entailed) |
| ✗ Fact 2: The father applied for denial of access. (Not-entailed) |
Appendix D Evaluation Prompts
Evaluation prompts for ARC , ARC, and ARC are described in Table 11, 12, 13 respectively. In each prompt, we ask the LLM to first generate a rationale before assigning a scoring decision following standard evaluation with LLMs Liu et al. (2023a).
| Prompt Given to GPT4-o for Fullset Evaluation |
| Task: |
| Evaluate how well a given summary covers a provided set of arguments. Assign a score from 1 to 4 based on the extent of coverage, provide a clear explanation for your rating, and output the result in a specified JSON format. |
| Instructions: |
| • Read the provided arguments and summary carefully. • Rate the extent to which the arguments are covered by the summary using the scale described in Table 8. • Format your evaluation as a JSON object with: – "explanation": A concise explanation of your rating. – "rating": The assigned score (1 to 4). |
| Example Output Format: |
| "explanation": "Place your explanation here", |
| "rating": "Place your rating here" |
| Input: |
| • Arguments: {reference_arguments} • Summary: {generated_summary} |
| Output: |
| Provide your evaluation in the specified JSON format. |
| Prompt Given to GPT4-o for Argument Role Evaluation |
| Task: |
| Determine whether a summary fully supports a given argument or omits/contradicts key information. |
| Instructions: |
| • Output 1 if the summary fully supports the argument without omissions or contradictions. • Output 0 if the summary fails to support the argument or contains contradictory or incorrect details (e.g., logical errors, entity mismatches). • Respond in a JSON object with: – "decision": Either 1 or 0. – "explanation": A brief justification, noting missing or conflicting content. |
| Input: |
| Argument: {argument} |
| Summary: {summary} |
| Output Format: |
| Respond only with a JSON object structured as: |
| "explanation": "<Brief reasoning for your decision>", |
| "decision": <0 or 1> |
| Note: Think critically before deciding. Do not include any extra text beyond the JSON output. |
| Prompt Given to GPT4-o for Atomic-Level Evaluation |
| Task Description: |
| Given an argument and a summary, evaluate whether the argument is supported by the summary and return a valid tuple in the specified format. |
| Explanation: |
| Provide a brief justification for your decision, identifying any missing, contradictory, or factually incorrect details. |
| Return Guidelines: |
| • (1, "supported"): The argument is fully supported by the summary. • (0, "missing"): The argument cannot be inferred from the summary. • (0, "not-factual"): The summary contradicts or misrepresents the argument. |
| Output Format: |
| Respond only with a JSON object, structured as: |
| "explanation": "<explanation placeholder>", |
| "decision": (1, "supported") or (0, "missing") or (0, "not-factual") |
| Input: |
| Argument: {argument} |
| Summary: {summary} |
| Note: Think critically before deciding. Do not generate any extra text beyond the JSON output. |
Appendix E Training and
We fine-tune DeBERTa, BigBirdRoBERTa, and LegalBERT using checkpoints obtained from the HuggingFace library Wolf et al. (2019). All models are trained using -fold cross-validation based on case-summary pairs for each model forming a total of case-summary pairs, resulting in argument-summary pairs for training and atomic fact-summary pairs for training . To handle class imbalance, we leverage weighted cross-entropy loss, where class weights are proportional to the label distribution across the dataset. This is particularly important for , where the label is significantly underrepresented compared to others. All models are trained for epochs using the Adam optimizer with an initial learning rate of . The maximum sequence length is set to for most models due to encoder constraints. However, for BigBirdRoBERTa, we expand the maximum length to to accommodate longer summaries and maintain coverage.
| Model | F1 Macro Score () | F1 Macro Score () |
| DeBERTa-base | 0.7138 | 0.7936 |
| DeBERTa-large | 0.7202 | 0.7936 |
| LegalBERT | 0.6850 | 0.7812 |
| BigBirdRoBERTa | 0.6629 | 0.7891 |
Table 14 shows that across all models, DeBERTa-Large obtained the best F1 macro average scores. Including encoders that can handle longer lengths didn’t improve the performance of prediction.
Appendix F ARC and ARC results
Figure 9 presents the results for both ARC and ARC across the CANLII and DRI datasets.
The ARC results indicate that both LlaMA-3.18B and Qwen2-7B achieve comparable levels of argument coverage from a holistic perspective. In contrast, and consistent with the ARC results presented in Section 5, Mistral-8B demonstrates slightly lower performance than both models. However, a key limitation of ARC is its insensitivity to fine-grained differences, as it primarily relies on coarse Likert-style scores that obscure nuanced variation across model outputs.
By contrast, ARC reveals that LlaMA-3.18B exhibits a higher number of perfectly covered roles compared to both Mistral-8B and Qwen2-7B. Nonetheless, due to its stricter definition of completeness at the role level, ARC underrepresents the strength of Qwen2-7B in capturing a greater number of atomic facts than Mistral-8B, as evidenced by both ARC and ARC.
Appendix G Correlation with source argument positions across full ARC scores
| Metric | Model | CANLII | DRI |
| ARC | LlaMA-3.1-8B | -0.137 | -0.152 |
| Mistral-8B | -0.042 | 0.250 | |
| Qwen2-7B | -0.187 | -0.174 | |
| ARC | LlaMA-3.1-8B | -0.200 | 0.202 |
| Mistral-8B | -0.244 | -0.065 | |
| Qwen2-7B | -0.216 | -0.124 | |
| ARC | LlaMA-3.1-8B | -0.230 | 0.323 |
| Mistral-8B | -0.369 | -0.055 | |
| Qwen2-7B | -0.301 | 0.041 |
Table 15 confirms that the LLM context window significantly impacts coverage across all scores (including ARC and ARC) in CANLII (), with a predominantly negative correlation. In contrast, its effect in DRI is not statistically significant, and in some cases (e.g., LlaMA-3.1-8B), it is slightly positive.
Appendix H Bias analysis for argument coverage without argument role normalization
While our normalization in computing corrects for frequency skew, it may understate coverage for dominant argument roles in the source with inherently high raw scores. Therefore, we also examine role-specific reporting bias by directly computing to offer a complete picture of the role bias analysis. As shown in Figure 10, results on CANLII confirm prior findings: LLMs tend to prioritize covering conclusion arguments over issue and reason roles, both in the controlled and non-controlled length settings. For DRI, removing frequency normalization reveals that ARC scores for roles such as background claim, own claim, and data are comparable to each other. This suggests that without normalization, the higher coverage of own claim may stem from its over-representation in the source documents, rather than a model-level preference.