### v2

# ARC: Argument Representation and Coverage Analysis for Zero-Shot Long Document Summarization with Instruction Following LLMs

Abstract
We introduce Argument Representation Coverage (ARC), a bottom-up evaluation framework that assesses how well summaries preserve salient arguments, a crucial issue in summarizing high-stakes domains such as law. ARC provides an interpretable lens by distinguishing between different information types to be covered and by separating omissions from factual errors. Using ARC, we evaluate summaries from eight open-weight LLMs in two domains where argument roles are central: long legal opinions and scientific articles. Our results show that while LLMs capture some salient roles, they frequently omit critical information, particularly when arguments are sparsely distributed across the input. Moreover, ARC uncovers systematic patterns—showing how context window positional bias and role-specific preferences shape argument coverage—providing actionable guidance for developing more complete and reliable summarization strategies.
1 Introduction
LLMs have achieved remarkable progress in abstractive summarization, producing summaries that are fluent, coherent, and often preferred by human evaluators in domains such as news Zhang et al. (2024); Liu et al. (2024b). This progress has shifted the focus of summarization research from how to generate fluent summaries to how to properly evaluate them. Evaluation is especially critical because, despite their fluency, LLMs frequently hallucinate or omit key content Huang et al. (2025); Kim et al. , undermining their reliability in high-stakes domains.
In this work, we focus on the evaluation of long-context summaries generated by instruction-following LLMs, focusing on whether they preserve salient argumentative content. In crucial domains such as law, the most essential information is conveyed through argument roles (the components of reasoning and main claims in the text). Preserving these roles is particularly challenging in long documents, where arguments are sparsely distributed across thousands of tokens Elaraby and Litman (2022). The ability of LLMs to identify and retain argument roles is therefore a crucial test of their utility for reliable summarization in high-stakes settings.
Why argument-based evaluation? Prior work has shown that explicitly modeling argument roles can improve summarization quality Fabbri et al. (2021b); Elaraby and Litman (2022); Elaraby et al. (2023). Building on this, argument-based evaluation allows us to address a fundamental research question: RQ1: Can instruction-following LLMs inherently identify and preserve argument roles without explicit supervision? Additonally, we see two key benefits of argument-based evaluation where existing metrics fall short: interpreting which types of information are missing or misrepresented in generated summaries, and guiding future alignment strategies for domain-specific summarization through a reproducible methodology. In this work, we focus on the first point by developing argument-based evaluation methods that make coverage interpretable.
We introduce Argument Representation Coverage (ARC), 111The full codebase, along with instructions for obtaining the data, is publicly available at https://github.com/EngSalem/ARCScore.git. a framework for evaluating how well LLM-generated summaries capture salient arguments. ARC leverages atomic fact decomposition of argument roles to measure coverage hierarchically. As illustrated in Figure 1 (decomposing roles to facts), all salient arguments are first decomposed into atomic facts to assess fact-level coverage. Each atomic fact is then compared against the generated summary and labeled as either supported, missing, or contradicted. These binary fact-level judgments are aggregated to compute role-level coverage (), providing fine-grained insight into how well each argument role is preserved. Finally, role-level scores are combined to produce summary-level coverage (), a holistic measure of how effectively a summary preserves salient argument roles. We apply ARC to study coverage in two domains where argument structure is essential for understanding the core of the document: long legal opinions and scientific articles.
Prior work shows that LLMs often exhibit positional biases—favoring content from the beginning or end of long documents Liu et al. (2024a); Ravaut et al. (2024); Wan et al. (2025). It is less clear, however, whether they also bias toward certain types of salient information that share the same structure (e.g., argument roles) during summarization. With ARC, we investigate additional two research questions: RQ2: How does the position of arguments in the source document affect their inclusion in summaries? and RQ3: Are certain argument roles disproportionately favored over others? To answer the latter, we leverage ARCrole to define a bias score that quantifies role-specific biases in saliency coverage.
Our analysis with ARC across eight open-weight LLMs reveals: instruction-following models still struggle with saliency coverage, particularly when argument roles are sparsely distributed, underscoring the need for further alignment; positional bias in the context window negatively impacts coverage; and bias analysis shows that LLMs exhibit systematic preferences for certain argument roles, particularly in long legal opinions.
| Dataset | # Docs | Input Length | Summary Length | % Roles in Input | % Roles in Summary |
| CANLII | |||||
| DRI | - |
2 Related Work
Information Saliency in LLMs’ Summaries. Content selection remains a core challenge in summarization. Trienes et al. (2025) found weak alignment between LLMs’ saliency preferences and human judgments. While LLMs can produce summaries preferred over human references in domains like news Zhang et al. (2024); Liu et al. (2024b), they still benefit from content planning. For example, Adams et al. (2023) showed that planning entity mentions improves the information density in GPT-4 generated summaries at the same summary length when compared to summaries generated without entity planning. In similar veins, Gantt et al. (2024); Walden et al. (2025) show that LLMs generate more focused and complete event-centered summaries when generation is explicitly conditioned on both the input document(s) and structured event representations. We extend this line of work by treating argument roles as a structured form of saliency and analyzing their preservation in LLM-generated summaries.
LLMs in Long-Document Summarization. LLMs face persistent issues when summarizing long texts, notably the U-shaped positional bias Liu et al. (2024a)—favoring content at the beginning and end while neglecting the middle Ravaut et al. (2024). This leads to degraded faithfulness in long-form outputs Wan et al. (2025). We expand this analysis by quantifying how positional bias affects the coverage of salient argumentative content.
Argument Mining in Abstractive Summarization. Incorporating argument structures into summarization has shown promise across domains, including dialogues Fabbri et al. (2021b), legal texts Xu et al. (2020, 2021); Elaraby and Litman (2022); Elaraby et al. (2023) and scientific documents Fisas et al. (2016). We build on this by assessing whether instruction-following LLMs can cover salient arguments without the external argument role information
| CANLII (Legal opinions) |
| Issue: Damage to both vehicles exceeded the insurance deductibles and both parties claim damages against each other. |
| Conclusion: Fault for this accident was attributed to the defendant and to the plaintiff. |
| Reason: Jurisdictional error is not to be equated with error of law. |
| DRI (Scientific documents) |
| Own Claim: Semi-Lagrangian contouring offers an elegant and effective means for surface tracking with advantages over competing methods. |
| Background Claim: Accurate modeling of human motion remains a challenging task. |
| Data: Animation is constrained due to hardware constraints. |
Evaluating Long-Form Summaries. Standard metrics often fall short in reflecting human preferences, especially for long documents Fabbri et al. (2021a); Krishna et al. (2023). To improve reliability, recent work has introduced unit-based metrics—such as Atomic Content Units (ACUs) Krishna et al. (2023) and structured factuality scores Min et al. (2023); Yang et al. (2024b)—to reduce subjectivity in evaluation. Extending this idea, we propose ARC, which uses argument roles as evaluation units and introduces subatomic granularity to assess fine-grained argument coverage.
3 Datasets
We employ two datasets that include both argument role annotations and reference summaries: CANLII Xu et al. (2021), representing the legal domain, and DR. INVENTOR (DRI) Fisas et al. (2016), representing the scientific domain. An overview of dataset statistics is presented in Table 1. Both datasets consist of long-form documents paired with long-form reference summaries that average words Krishna et al. (2023). More analysis and examples are given in Appendix A.
3.1 Legal Opinions: CANLII
The CANLII dataset consists of legal cases annotated at the sentence level for argument roles. Notably, only of the input text is labeled with argument roles, yet these argumentative sentences account for of the reference summaries (Table 1). This substantial mismatch highlights a haystack-like challenge: models must accurately identify and prioritize the sparse yet highly salient argumentative content when generating summaries.
Argument roles in CANLII are annotated using the IRC scheme Xu et al. (2021), which categorizes roles into three types: Issues: Legal questions raised in the case. Reasons: Justifications provided for judicial decisions. Conclusions: Final rulings addressing the identified issues. These annotations enable a fine-grained evaluation of argument coverage in generated summaries. Examples of IRCs are shown in Table 2.
3.2 Scientific Articles: DRI
The DRI dataset consists of computer graphics articles, each annotated at the sentence level for rhetorical roles and paired with three human-written summaries. Notably, these rhetorical roles are not necessarily argumentative. To address this limitation, the extended version of the dataset, SCI-ARG Lauscher et al. (2018), enriches the DRI annotations by incorporating argument role annotations and their relations. These annotations follow a modification of the -argument roles described in Toulmin model Toulmin (2003), by reducing them into three types: Own Claim: Sentences that directly support the author’s central argument. Background Claim: Sentences that reference prior research or established domain knowledge. Data: Empirical evidence that supports or refutes claims, such as experimental results or literature citations. An example of each role is shown in Table 2.
Since the argument annotations are span-based, we map them back to complete sentences using lexical matching assigning the sentence with argument role spans if of its words falls within the sentence boundaries. A motivating feature behind selecting this corpus for our analysis is the sentence-level annotation for relevance scores on a Likert scale , which indicates the degree of relevance to the summary. A relevance score of signifies that the sentence is "relevant to the summary," while a indicates it is "very relevant to the summary." In our evaluation, we focus on argument role coverage for sentences with argument roles and a Likert score of (indicating high relevant argument roles to the summary). Table 1 shows that unlike legal opinions, where argument roles are sparsely distributed, scientific articles contain argumentative content throughout the document, posing a challenge of selectivity rather than retrieval. In DRI, sentences that contain at least 1 argument role account for of the input text (shown in Table 1). Although the dataset does not provide gold-standard summaries annotated for argument roles, we analyze the sentences with a Likert score of based on their argument role annotation. Among these sentences, contain at least one argument role, reinforcing the strong connection between argument roles and summarization relevance in this domain.
4 The ARC Framework
4.1 Overview and Notations
Decomposing Roles to Facts. For each argument role , where is the set of salient argument roles (roles that are essential to be included in the summary), we follow the decomposition algorithm of Yang et al. (2024b). Specifically, we employ a strong LLM (GPT4-o) to decompose into a set of atomic facts . These facts are further filtered using an entailment check against to eliminate over-generated facts. 222Decomposition algorithm in Appendix B.
Role-Level Coverage. Given a generated summary , we define the indicator function:
The role-level coverage for is then:
Summary-Level Coverage. The overall coverage score, ARCscore, across all roles is computed as aggregation of across :
4.2 Evaluating Factual Units ()
The main goal of the ARC framework is to provide an interpretable evaluation of coverage in generated summaries. To do so, the function distinguishes between two error types: missing facts (information omitted) and incorrectly covered facts (information misrepresented). 333Examples of atomic facts are included in Appendix C. Although both reduce completeness, separating them reveals whether errors stem from failing to prioritize salient content or from factual mistakes during generation. We implement using an LLM judge. 444See Appendix D for the evaluation prompt. We employ zero-shot LLMs, which have shown strong correlation with human judgments in summary evaluation Liu et al. (2023). Given an atomic fact and a summary, we prompt the LLM to return a pair , where indicates whether the fact is supported (covered) or not supported (missing or incorrect), and provides a categorical interpretation of the error type (no error, missing, non-factual).
4.3 Benchmarking ARCscore
To evaluate the effectiveness of ARCscore in capturing holistic summary coverage, we benchmark it against existing automatic metrics using expert-annotated data.
Expert-Annotated Data. Human-annotated datasets for fine-grained coverage evaluation are extremely scarce. To our knowledge, the only available resource is the dataset of Elaraby et al. (2024), which includes legal opinion summaries annotated for salient argument coverage on a -point Likert scale. 555See Appendix E for scale definitions. Two legal experts independently annotated the data, achieving a moderate agreement of (quadratic weighted). To improve the robustness of the annotations, we filtered out pairs with point of disagreement, resulting in article–summary pairs with improved agreement (). We use the expert average as the gold standard to compare against automatic metrics.
Automatic Metrics for Holistic Coverage. We benchmark ARC against a suite of strong automatic metrics. For all baselines, we treat the set of salient arguments as the hypothesis and the generated summary as the premise/reference. We first include the standard metrics reported in Elaraby et al. (2024): ROUGE-1, ROUGE-2, ROUGE-L, and BERTScore. Next, we add entailment-based baselines from Laban et al. (2022), where SummaC computes entailment between hypothesis and reference at both sentence-level and document-level granularity. We include both (zeroshot zs and convolution conv versions of SummaC). Finally, we incorporate an LLM-judge baseline: FactScore Min et al. (2023), which—similar to ARCscore—relies on decomposition for fact-level evaluation. We adapt FactScore for coverage by treating the generated summary as the knowledge base and the salient arguments as hypotheses to be tested. For fairness, both ARCscore and FactScore use GPT4-o as the evaluator. We report correlation with expert averages using Kendall’s and Pearson’s .
| Metric | Expert Avg. | Interpretability | ||
| Info Types | Errors | |||
| ROUGE-1 | ✗ | ✗ | ||
| ROUGE-2 | ✗ | ✗ | ||
| ROUGE-L | ✗ | ✗ | ||
| BERTScore | ✗ | ✗ | ||
| SummaCZS (sent) | ✗ | ✗ | ||
| SummaCZS (doc) | ✗ | ✗ | ||
| SummaCconv (sent) | ✗ | ✗ | ||
| SummaCconv (doc) | ✗ | ✗ | ||
| FactScore (GPT4-o) | ✗ | ✗ | ||
| ARCscore (GPT4-o) | ✓ | ✓ | ||
| Judge Model | ||
| Proprietary models | ||
| GPT-4o | 0.465 | 0.593 |
| GPT-4o-mini | 0.446 | 0.582 |
| Open-weight Instruction-following models | ||
| Qwen-2.5-7B-Instruct | 0.329 | 0.480 |
| Qwen-2.5-14B-Instruct | 0.454 | 0.601 |
| Llama-3.1-8B-Instruct | 0.463 | 0.610 |
| Mistral-8B-Instruct | 0.443 | 0.599 |
| Open-weight Reasoning models | ||
| QwQ-32B-AwQ | 0.437 | 0.596 |
| DeepSeek-R1-Distill-Qwen-14B | 0.509 | 0.638 |
Table 3 shows that among automatic metics, ARCscore achieves the highest correlation with expert judgments,666See Appendix F for correlations with each expert. outperforming standard evaluation metrics and LLM-based decomposition-based metrics such as FactScore. Crucially, ARCscore improves the holistic view of coverage over baselines, while also providing interpretability that existing metrics lack. Unlike FactScore, which provides only fact-level coverage, ARC enables multi-level analysis: role-level coverage via ARCrole, and error-type analysis distinguishing between missing and incorrect facts. This decomposition allows us to diagnose not only how much coverage is achieved, but also which roles are underrepresented and why errors occur, making ARC a more effective diagnostic tool for granular coverage evaluation.
Alternative LLM-judges. To make ARC more scalable and accessible, we explore alternatives to GPT-4o for fact-level evaluation that maintain high correlation with expert ratings while reducing inference cost and avoiding reliance on proprietary APIs that may be deprecated. We focus on two categories of open-weight models: instruction-following models. We evaluate smaller open-weight models that fit within our compute and memory budgets, including Qwen-2.5-Instruct (7B and 14B), Llama-3.1-8B-Instruct, and Mistral-8B-Instruct. Reasoning models. We further examine open-weight reasoning models, specifically QwQ-32B-AwQ and DeepSeek-R1-Distill-Qwen-14B, which have demonstrated competitive performance with proprietary instruction-following and reasoning models on complex tasks, including verification. Table 4 shows that DeepSeek-R1-Distill-Qwen-14B achieves the strongest overall correlation, surpassing even proprietary models such as GPT-4o. Among open-weight instruction-following models, Llama-3.1-8B-Instruct attains the highest correlation, outperforming GPT-4o-mini. We employ both proprietary (GPT-4o) and open-weight judge (DeepSeek-R1-Distill-Qwen-14B) to compute ARCscore. To balance cost and scalability, GPT-4o is applied to a representative subset of CANLII (CANLII100) cases and the full DRI corpus, while DeepSeek-R1-Distill-Qwen-14B is used for both CANLII100 and the full CANLII ( articles). Our goal is to ensure that the sensitivity to the underlying judge (i.e., disagreements at the fact level) does not alter the overall coverage trends observed across models.
| Model | CANLII100 | CANLII | DRI | ||
| GPT4-o | DeepSeek | DeepSeek | GPT4-o | DeepSeek | |
| Reference | 0.986 | 0.977 | 0.982 | 0.838 | 0.813 |
| Qwen-2.5-14B | 0.677 | 0.707 | 0.722 | 0.772 | 0.797 |
| Qwen-2.5-7B | 0.676 | 0.691 | 0.725 | 0.799 | 0.806 |
| Qwen-2.5-3B | 0.648 | 0.702 | 0.710 | 0.730 | 0.759 |
| Qwen-2.5-1.5B | 0.472 | 0.512 | 0.539 | 0.668 | 0.628 |
| Mistral-8B | 0.554 | 0.609 | 0.608 | 0.669 | 0.697 |
| LLaMA-3.1-8B | 0.641 | 0.642 | 0.678 | 0.749 | 0.793 |
| LLaMA-3.2-3B | 0.573 | 0.603 | 0.639 | 0.705 | 0.740 |
| LLaMA-3.2-1B | 0.451 | 0.530 | 0.545 | 0.674 | 0.631 |
5 Analyzing Summary Coverage with ARC
In this section, we use ARC to evaluate summaries produced by long-context LLMs on the CANLII and DRI datasets.
5.1 Obtaining Generated Summaries
We evaluate eight long-context open-weight LLMs from the LLaMA Grattafiori et al. (2024), Mistral Jiang et al. (2024) and Qwen Yang et al. (2024a); Team (2024) families that meet our computational constraints. Specifically, we include LLaMA-3.1-8B-Instruct, LLaMA-3.2-1B and 3B, Qwen-2.5-Instruct-1.5B, 3B, 7B, 14B, and Mistral-8B. We vary model sizes to examine whether scaling improves coverage performance. Inference is conducted using VLLM Kwon et al. (2023) for efficiency and scalability, with Qwen models further extended to tokens via RoPE scaling (factor ).
Summaries are generated using greedy decoding (), capped at tokens to ensure fair long-form generation. Each document is prompted with the following instruction: "Read the following text and summarize it: {input document}. Summarize in {reference summary word length} words. Summary:" We deliberately use a generic summarization prompt to evaluate LLMs without giving them any prior indication of which information is salient.777See Appendix G for experiments with an argument-aware prompt. This setup mirrors the human annotation process, where experts also lacked explicit guidance on saliency when producing reference summaries.
| CANLII100 | CANLII | DRI | ||||||||
| Model | MF | FE | MF | FE | MF | FE | MF | FE | MF | FE |
| Qwen-2.5-14B | 35.7 | 2.0 | 31.4 | 3.2 | 28.7 | 3.1 | 36.5 | 0.2 | 36.7 | 0.5 |
| Qwen-2.5-7B | 35.5 | 2.0 | 31.2 | 3.6 | 27.8 | 3.2 | 33.5 | 0.5 | 34.7 | 1.0 |
| Qwen-2.5-3B | 36.2 | 2.8 | 30.5 | 3.4 | 29.5 | 3.6 | 36.6 | 0.8 | 36.6 | 0.7 |
| Qwen-2.5-1.5B | 56.3 | 3.6 | 49.1 | 5.5 | 46.9 | 4.9 | 49.9 | 0.4 | 47.8 | 1.1 |
| Mistral-8B | 46.8 | 3.0 | 42.1 | 3.1 | 40.4 | 3.7 | 43.3 | 0.4 | 44.0 | 1.3 |
| LLaMA-3.1-8B | 40.1 | 2.1 | 36.9 | 3.6 | 33.1 | 3.3 | 41.1 | 0.3 | 41.7 | 0.8 |
| LLaMA-3.2-3B | 45.4 | 3.0 | 40.1 | 4.0 | 36.5 | 4.2 | 43.8 | 0.3 | 43.3 | 1.3 |
| LLaMA-3.2-1B | 56.3 | 4.6 | 49.1 | 5.5 | 45.1 | 5.7 | 49.9 | 0.7 | 50.0 | 1.3 |
The target length is dynamically matched to the reference summary for comparability. For DRI, the target length is fixed to the longest reference summary to encourage maximal argument retention.
5.2 RQ1: Can LLM summaries cover salient arguments without explicit supervision?
We compute ARCscore for all model-generated summaries described in Section 5.1, as well as for human reference summaries on both the CANLII and DRI benchmarks. For CANLII, we report results on both the smaller -article subset (CANLII100) scored with GPT-4o and the full set (CANLII) scored with DeepSeek-R1-Distill-Qwen14B, thereby testing the robustness of ARCscore across different judges and verifying that it assigns near-perfect scores to expert-written summaries. For DRI, because role-level annotations in summaries are unavailable, the goal is to assess how well human references capture salient argument roles compared to LLM-generated outputs.
Table 5 shows that argument coverage remains incomplete across all models and domains, indicating that LLMs could still benefit from argument-aware supervision. Although the two judges yield slightly different ARCscore values, the overall ranking of models is consistent. 888For CANLII100, Kendall’s (between the judges scores); for DRI, ; both . Two clear patterns emerge from the results. Larger models generally achieve higher coverage. Within the Qwen-2.5 family, performance improves when scaling from B to B and B across both judges. A similar trend holds for LLaMA-3.2, where the B model surpasses the B variant. 999This scaling effect is examined only for models with B–B parameters. Coverage is more challenging in legal opinions. On DRI, the strongest model (Qwen-2.5-7B) achieves , closely approaching the human reference () under both judges. By contrast, on CANLII100 and the full CANLII, the best-performing models plateau at and , far below the human reference ( and , respectively). This disparity underscores the difficulty of preserving salient argumentative content in legal texts, where arguments (both salient and non-salient) are sparsely distributed across lengthy contexts.
Fact-level Analysis. We analyze fact-level decisions (missing versus non-factual) to determine whether coverage errors arise primarily from omissions of key information or from factual inaccuracies. For each model, we compute the proportion in of missing facts and factual errors across all generated summaries. As shown in Table 6, missing facts are the dominant source of error across both datasets, judges, and models, indicating that salient information is more often omitted than misrepresented. While factual inconsistencies do occur, they are consistently less frequent. These findings suggest that, beyond hallucination, the central challenge in summarization is achieving comprehensive coverage of salient content.
5.3 RQ2: Do argument positions in the source affect their coverage in summaries?
Following Ravaut et al. (2024), we start by analyzing the positions where included LLMs look at in its context window. We leverage the lexical greedy approach for source sentences identification Ravaut et al. (2024); Adams et al. (2023) by iteratively adding sentences in the source that maximizes ROUGE-1 score until there is no further improvement. 101010Appendix H contains the full algorithm. We analyze the source sentence indices by their argument role annotations. Figure 2 shows a clear U-shaped context window bias across all models, most pronounced in CANLII. Argument role analysis indicates that source positions are heavily shaped by this pattern—problematic for CANLII, where reference summaries do not follow fixed positional trends. In contrast, DRI references align more closely with the LLM bias distribution.
| CANLII100 | CANLII | DRI | |||
| Model | |||||
| Qwen-2.5-14B | -0.144 | -0.131 | -0.146 | 0.119 | -0.044 |
| Qwen-2.5-7B | -0.301 | -0.237 | -0.214 | -0.223 | 0.007 |
| Qwen-2.5-3B | -0.232 | -0.112 | -0.136 | 0.006 | 0.039 |
| Qwen-2.5-1.5B | -0.230 | -0.170 | -0.163 | 0.074 | -0.010 |
| Mistral-8B | -0.369 | -0.244 | -0.193 | -0.055 | 0.012 |
| LLaMA-3.1-8B | -0.230 | -0.088 | -0.153 | 0.129 | 0.166 |
| LLaMA-3.2-3B | -0.216 | -0.113 | -0.190 | -0.091 | -0.031 |
| LLaMA-3.2-1B | -0.171 | -0.087 | -0.198 | -0.020 | -0.031 |
We analyze how the position of salient arguments affects coverage as measured by . For CANLII and CANLII100, we apply greedy sentence selection while restricting the pool to annotated arguments that appear in both the reference summary and the input document. This reduces computational cost and ensures that only arguments included in the reference summary are mapped back to the input. For DRI, we directly identify the positions of arguments in the input document with a relevance score of . Following Ravaut et al. (2024), we compute the mean relative position of salient arguments and measure the Pearson correlation between this mean position and .
Table 7 shows a consistent negative correlation in CANLII (both CANLII100 with GPT4-o and DeepSeek–R1, as well as the full CANLII set with DeepSeek–R1), with several models reaching significance (). This confirms that LLM context windows systematically bias coverage in long legal cases. By contrast, correlations in DRI are weaker and largely non-significant, with some models (e.g., LLaMA-3.1-8B) even showing slight positive trends under DeepSeek–R1. These results align with Figure 2: in CANLII, reference summaries diverge from the strong U-shaped bias present in model outputs, whereas in DRI, reference arguments more closely mirror the positional distribution induced by the LLM context window.
5.4 RQ3: Are argument roles disproportionately covered?
We propose as a role-specific bias score, designed so that higher values indicate stronger bias (i.e., poorer representation of a role). The intuition is that if a role is perfectly represented in the summary, then , and the bias should vanish (). Conversely, lower coverage implies a larger gap from perfect representation, reflected in a higher . Formally, we define:
where is the frequency of role in the source document , and is the total number of arguments in . The normalization term downweights roles that are overrepresented in the source, ensuring that captures true disparities in coverage rather than frequency effects.
To further reduce confounds from argument length and position—especially in CANLII,111111This analysis uses CANLII100 with the GPT4-o judge, as results in RQ1 and RQ2 show consistent trends across judges and dataset splits. where longer roles and arguments positioned in the middle of the document can bias coverage—we compute the bias score within length-controlled groups, allowing a variation in word count. In addition, to control for positional bias effect on coverage, we restrict the positional analysis to cases where at least 80% of arguments fall within either the first or the last 20% of the document. Positions are computed using the relative positions of the indices of argumentative sentences from the beginning of the document. These constraints help ensure that the observed effects on role coverage are not confounded by argument length or positional bias.
Figure 3121212Appendix I reports results without frequency normalization to rule out denominator inflation effects. shows that, for CANLII100, is consistently lowest for conclusions, indicating stronger coverage of this role across both the all-argument analysis and the position-controlled setting. Controlling for role length explains part of the variation across roles, but the relative advantage of conclusions persists. In DRI, by contrast, bias scores are generally lower than in CANLII100, consistent with our earlier finding that LLMs capture more salient arguments in scientific summaries. Nonetheless, background claims tend to be covered less effectively, even when length is controlled. Across both domains, model size further influences role-level disparities: for instance, models with B parameters consistently exhibit the highest , reflecting weaker and less balanced argument coverage.
6 Conclusion and Future Work
We introduced ARC, a hierarchical evaluation framework for assessing how well LLM-generated summaries preserve salient argumentative content. ARC provides a principled and interpretable diagnostic tool for evaluating structured argument coverage in long-context summarization. Its hierarchical structure not only enables interpretability—revealing which roles are preserved and whether errors arise from omissions or factual inaccuracies—but also achieves higher correlation in holistic evaluation compared to lexical, semantic, entailment, and decomposition-based baselines. Applying ARC to legal and scientific domains uncovered two consistent limitations of current LLMs: positional bias, where the characteristic U-shaped context window negatively affects coverage, and role bias, where conclusions are favored over other roles such as issues and reasons. Future work can extend this framework by incorporating explicit argument structures into training or prompting, and by leveraging ARC’s interpretable outputs to guide targeted improvements in summarization—particularly in high-stakes domains.
Limitations
While the ARC framework enables a comprehensive, multi-level evaluation of argument coverage, several limitations remain that suggest promising directions for future work.
Limited Benchmarks. Evaluation is constrained by the scarcity of benchmarks explicitly designed for argument coverage. Existing datasets provide limited annotation granularity, especially below the argument-unit level. To our knowledge, only the two benchmarks used here include both salience annotations and explicit argument roles. The small size of DRI (40 documents) also limits generalizability, motivating larger, rigorously annotated corpora—e.g., debates or financial texts—where arguments are central.
Decomposition Sensitivity. Our decomposition into atomic facts relies on the Yang et al. (2024b) algorithm, LLM-based prompting, and entailment filtering, which may introduce errors. While this decomposition proved more effective than that in FactScore Min et al. (2023), it is worth noting that FactScore relied on human-authored examples from biographical texts, which might have contributed to their lower correlations. Future extensions of ARC should consider incorporating expert-written atomic facts to further improve consistency and balance.
Balancing Precision and Recall. ARC currently emphasizes recall—whether salient argument-role facts are preserved—similar to FactScore. While critical for high-stakes domains, this ignores precision, i.e., over-inclusion of non-salient content. Future work should jointly assess both dimensions, for instance via a harmonic mean, for a fuller view of coverage quality. 131313We additionally conduct a precision–recall analysis on a subset of CANLII summaries; details and results are reported in Appendix J.
Dependence on Gold Annotations. Our analysis assumes access to gold salience labels. Though this allows conclusive evaluation, future research should explore end-to-end systems that integrate salience detection, fact decomposition, and coverage assessment within a unified framework.
Ethics Statement
Our study complies with the ACL Ethics Policy. We primarily evaluate academically available datasets designed explicitly for research purposes,which we obtained through license agreement with the authors of both datasets, thus minimizing privacy risks. Additionally, our work acknowledges potential biases and inaccuracies inherent to LLM-generated outputs, including misrepresentation or omission of critical information from summaries, which could have significant implications in high-stakes domains such as law and science. Researchers and practitioners utilizing our framework should exercise caution and validate results carefully before applying these models in sensitive or consequential decision-making contexts.
Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. 2040490 and by Amazon. This research was supported in part by the University of Pittsburgh Center for Research Computing through the resources provided. This research was partially supported by the CS50 fellowship awarded by the department of computer science at the University of Pittsburgh. We want to thank the members of the Pitt PETAL group, Pitt NLP group, and anonymous reviewers for their valuable comments in improving this work.
Appendix A Extended analysis on included datasets
A.1 Examples from included datasets
Table 8 presents an excerpt from a legal opinion in the CANLII dataset, with arguments highlighted in both the input and the reference summary. Table 9 provides an excerpt from a scientific article in the DRI dataset, with highlighted arguments in the input. Although the documents are truncated for space, the examples clearly illustrate a key distinction: in CANLII, arguments constitute a smaller fraction of the input, whereas in DRI, the input is densely populated with argumentative content.
| Input Article (Truncated) |
| Q.B. A.D. 1987 No. CS 1159 J.C.R., Regina. Applicants seek to quash a search warrant issued by a justice of the peace. The respondent, a justice of the peace, issued a search warrant to search a dwelling house for weapons allegedly used in an attempted armed robbery. The applicants claim the warrant was unlawfully issued without proper grounds. Specifically, the sworn information relied solely on hearsay from an unidentified informant, lacking corroborating details. The applicants argue that no reasonable or probable grounds were disclosed to believe the weapons would be found at the searched location. They highlight that the informant’s reliability was not established, nor was there an oath affirming the informant’s credibility. The search warrant was issued under Section 443(1)(b) of the Criminal Code, which allows a justice to issue a warrant if reasonable grounds exist to believe evidence of an offence will be found. The court explains that on applications to quash a warrant, the reviewing judge cannot substitute their opinion for that of the justice of the peace. Instead, the judge must simply determine whether any evidence existed upon which the justice could be satisfied on reasonable grounds. Reliance on confidential informants is permitted, even if detailed particulars are absent, provided sufficient basis exists for reliability. Past cases (e.g., Re Lubell, Re Dodge) have accepted similar levels of disclosure to protect informant anonymity. The court notes that substantial compliance with Section 443 is sufficient; perfection in drafting is not required. Given practical constraints faced by peace officers preparing information, reasonable latitude must be given in interpreting the sworn information. The judge concludes that although more information could have been provided, there was sufficient evidence upon which the justice could reasonably issue the warrant. Accordingly, the respondent acted within her jurisdiction, and the application to quash the warrant is dismissed. |
| Reference Summary |
| Warrant issued to search a dwelling house for weapons allegedly used in an attempted armed robbery. The affidavit in support referred to an unknown informant. Judge applied the test that the justice of the peace ‘must be satisfied on reasonable grounds.’ Substantial compliance found and warrant upheld. |
| Input Article (Truncated) |
| Our method maintains an explicit polygonal mesh that defines the surface, and an octree data structure that provides both a spatial index for the mesh and a means for efficiently approximating the signed distance to the surface. At each timestep, a new surface is constructed by extracting the zero set of an advected signed-distance function. Semi-Lagrangian backward path tracing is used to advect the signed-distance function. One of the primary advantages of this formulation is that it enables tracking of surface characteristics, such as color or texture coordinates, at negligible additional cost. We include several examples demonstrating that the method can be effectively used as part of a fluid simulation to animate complex and interesting fluid behaviors. The fundamental problem of tracking a surface as it is advected by some velocity field arises frequently in applications such as surface reconstruction, image segmentation, and fluid simulation. Unfortunately, the naive approach of simply advecting the vertices of a polygonal mesh quickly encounters problems such as tangling and self-intersection. Instead, a family of methods, known as level-set methods, has been developed for surface tracking. These methods represent the surface implicitly as the zero set of a scalar field defined over the domain. Level-set methods avoid dealing with topological changes but require high-order conservation law solvers. In contrast, our method constructs a surface directly using semi-Lagrangian contouring without solving PDEs, preserving surface detail efficiently. Using adaptive octree data structures, we can efficiently and reliably construct the new surface and corresponding signed-distance function. This allows tracking surface properties such as color or texture coordinates directly on the polygonal mesh during advection, enabling realistic animation of complex fluids. Prior methods often suffered from volume loss and smoothing artifacts, particularly in underresolved, high-curvature regions. By using an explicit surface representation, we compute exact distances near the mesh and avoid substantial interpolation errors. … Finally, the method produces detailed, flicker-free animations of fluid behavior, demonstrating significant advantages over traditional level-set and particle-based approaches. |
| Reference Summary |
| This article presents a semi-Lagrangian surface tracking method that explicitly represents the surface as a set of polygons. The new surface and corresponding signed-distance function can be efficiently and reliably constructed using adaptive octree data structures. One of the primary advantages of this method is that it enables tracking surface characteristics, such as color or texture coordinates, or even simulation variables, accurately at negligible additional cost. These properties can be easily stored directly on the polygonal mesh and efficiently mapped onto the new surface during semi-Lagrangian advection. At each timestep, a new surface is constructed by extracting the zero set of an advected signed-distance function. The explicit representation provides advantages on computing exact signed-distance values near the mesh and storing properties on mesh vertices. It also facilitates other common operations developed for manipulating and rendering triangle meshes. To avoid the topological difficulties of directly updating an explicit surface representation, the surface is updated in time through an implicit representation. The implicit representation is then used to construct a new mesh and extracted using a contouring algorithm. For its simplicity, robustness, and speed, marching-cubes method is used for contouring. After the triangle mesh has been extracted, true distance values are assigned to the vertices of octree. This process is known as redistancing, which comprises three steps: coarsen the octree; compute exact distances at vertices; run a fast marching method over the remaining vertices. Finally, this method is able to produce detailed, flicker-free animations of complex fluid motions. |
A.2 Distribution of arguments across the input
Figure 4 illustrates the distribution of argument roles across the source documents. In CANLII, Conclusion statements predominantly appear toward the end of the document, while Issue statements are concentrated near the beginning. In DRI, Background claims are more frequent at the start of the document, which aligns with the conventional structure of scientific writing where literature reviews—typically containing claims from prior work—are introduced early on.
A.3 Distribution of salient arguments
Figure 5 presents the distribution of argument roles in CANLII reference summaries and in DRI sentences annotated with a relevance Likert score of 5 (indicating very high likelihood of inclusion in a summary). In CANLII, the distribution of argument roles is relatively balanced across categories, whereas in DRI, own claims—statements made directly by the authors—dominate the content.
A.4 Rhetorical roles per relevance to summary Likert score
To better understand the relationship between rhetorical structure and relevance to the summary, we compute the percentage of each rhetorical role across Likert-rated sentences in the DRI corpus. As shown in Figure 6, non-argumentative content dominates among sentences rated as totally irrelevant to the summary (Likert score 1). However, as the perceived relevance increases, argumentative content becomes more prominent, with Own Claim consistently emerging as the most frequent rhetorical role across all higher-quality categories. This trend highlights a clear shift toward structured argumentative writing in more relevant argument roles.
Appendix B Fact decomposition algorithm
Algorithm 1 outlines the decomposition process for an arbitrary argument , performed via prompting GPT-4o.
| Prompt Given to GPT4-o for Argument Decomposition |
| Task: |
| Extract a set of atomic facts—statements that can be directly inferred from the argument without interpretation, assumptions, or redundancy. |
| Guidelines: |
| • Extract only explicitly stated atomic facts. • Do not repeat facts or infer from external knowledge. • Maintain granularity: each fact should be minimal yet complete. • Output a valid Dictionary object where each key is "fact1", "fact2", etc., and the values are the corresponding atomic facts. • No additional text or formatting; dictionary object only. • Each argument must yield at least one atomic fact. |
| Example Output Format: |
| "fact1": "First atomic fact", |
| "fact2": "Second atomic fact", |
| "fact3": "Third atomic fact" |
| Input: |
| {argument} |
| Output: |
| (Dictionary object only) |
| Argument (Issue): |
| FIAT. The father applied to have the mother cited for contempt for denial of access. |
| ✓ Fact 1: The father applied to have the mother cited for contempt. (Entailed) |
| ✗ Fact 2: The father applied for denial of access. (Not-entailed) |
Appendix C Atomic fact error examples
Table 12 presents representative examples from the DRI dataset, illustrating model decisions on atomic facts as evaluated by GPT-4o.
| Example 1 – Non-factual |
| Summary: The paper presents an efficient fluid simulation method using Laplacian eigenfunctions as a velocity basis, which avoids the high computational complexity of traditional Eulerian and Lagrangian approaches. This method is particularly advantageous for incompressible fluid dynamics, offering a principled way to reduce the dimensionality of the Navier–Stokes equations. The authors argue that Laplacian eigenfunctions provide a natural correspondence with spatial scales of vorticity, allowing precise control over turbulent energy at different scales. They show that for simple domains, these eigenfunctions have closed-form expressions, enabling fast and exact evaluation of the velocity field. The method is formulated using Galerkin projection onto Laplacian eigenfunctions, which are divergence-free and satisfy free-slip boundary conditions. For complex domains, a discrete exterior calculus (DEC) approach is used to compute the basis fields. The paper highlights the benefits of this method, including its ability to simulate fluid motion interactively, its low memory and precomputation requirements, and its flexibility in controlling viscosity and external forces. However, it also acknowledges limitations, such as stair-step artifacts on irregular boundaries and the need for additional methods to handle moving obstacles. The authors demonstrate that the method can be applied to various domains, including simple rectangles and complex 3D cavities, and discuss its potential for expressive control and optimization in fluid dynamics. The paper concludes by noting that while the method is particularly suited for gaseous phenomena, it may require further development for liquid simulations. |
| Fact: Our basis has a natural correspondence with spatial scales of vorticity that is lacking in [Treuille et al. 2006]. Decision: Non-factual Explanation: The summary correctly states that the basis corresponds to vorticity scales but omits the comparative qualifier “lacking in [Treuille et al. 2006],” misleadingly implying a novel rather than comparative contribution. |
| Example 2 – Missing Fact |
| Summary: Pose space deformation (PSD) is a technique that unifies and improves upon both shape interpolation and skeleton-driven deformation methods. It represents deformations as mappings from a pose space to displacements in the object’s local coordinate frames. PSD allows direct manipulation of desired shapes and maintains the performance benefits of traditional methods. Key applications include facial and body deformation in entertainment, telepresence, and gaming. The limitations of shape interpolation and skeleton subspace deformation (SSD) motivate the development of PSD. Shape interpolation lacks smoothness and requires frequent sculpting of new key shapes. SSD, while versatile, suffers from the "collapsing elbow" problem and is difficult to control. PSD addresses these issues by interpolating deformations in a pose space, enabling smooth and controllable deformations. PSD requires scattered data interpolation in high-dimensional spaces. Gaussian radial basis functions are used due to their well-behaved nature and ease of implementation. The algorithm is simple, general, and allows real-time synthesis, making it suitable for high-resolution models. Facial animation can benefit from PSD by allowing sculpting of intermediate expressions and multidimensional pose spaces. PSD can handle various deformation scenarios uniformly, from simple joints to complex secondary animations. The setup and synthesis costs are minimal, ensuring real-time performance. While anatomically based models offer higher quality, PSD provides a practical alternative for real-time applications and fanciful creature designs. |
| Fact: The synthesis cost is only slightly more than that of shape interpolation. Decision: Missing Fact Explanation: The summary omits the comparative detail that PSD’s synthesis cost is slightly higher than that of shape interpolation, overstating the efficiency of the proposed method. |
Appendix D Evaluation prompt
Evaluation prompt for fact-level coverage ARC is described in Table 13. We ask the LLM to generate a rationale before assigning a decision.
| Prompt Given to GPT4-o for Atomic-Level Evaluation |
| Task Description: |
| Given an argument and a summary, evaluate whether the argument is supported by the summary and return a valid tuple in the specified format. |
| Explanation: |
| Provide a brief justification for your decision, identifying any missing, contradictory, or factually incorrect details. |
| Return Guidelines: |
| • (1, "supported"): The argument is fully supported by the summary. • (0, "missing"): The argument cannot be inferred from the summary. • (0, "not-factual"): The summary contradicts or misrepresents the argument. |
| Output Format: |
| Respond only with a JSON object, structured as: |
| "explanation": "<explanation placeholder>", |
| "decision": (1, "supported") or (0, "missing") or (0, "not-factual") |
| Input: |
| Argument: {argument} |
| Summary: {summary} |
| Note: Think critically before deciding. Do not generate any extra text beyond the JSON output. |
Appendix E Likert scores based on human evaluation
Table 14 shows the Likert scale from to definitions.
| Rating | Explanation of the Generated Summary Rating Scale |
| 1 | No arguments covered: The generated summary did not cover the highlighted arguments in the reference summary or covered them only inadequately. |
| 2 | Few arguments covered: The generated summary adequately covered only a limited number of the highlighted arguments in the reference summary. |
| 3 | Most arguments covered: The generated summary adequately covered most of the arguments highlighted in the reference summary. |
| 4 | All arguments covered: The generated summary adequately covered all the highlighted arguments in the reference summary. |
Appendix F Full expert correlation
Table 15 reports correlations with individual experts and their averaged ratings across both judges. ARCscore (with DeepSeek-R1-Distill-Qwen14B) achieves the strongest overall correlation with expert judgments, surpassing all baselines. ARCscore (with GPT4-o) yields the second-highest overall correlation, outperforming all metrics except SummaCZS (sentence-level) on Expert 2. These results suggest that the decomposition in ARCscore not only enables fine-grained and interpretable analyses across error types and argument-role distinctions, but also preserves—and even improves upon—holistic coverage evaluation compared to strong automatic baselines.
| Metric | Expert 1 | Expert 2 | Average | |||
| ROUGE-1 | 0.378 | 0.520 | 0.347 | 0.456 | 0.391 | 0.539 |
| ROUGE-2 | 0.363 | 0.441 | 0.276 | 0.419 | 0.336 | 0.475 |
| ROUGE-L | 0.310 | 0.417 | 0.337 | 0.450 | 0.345 | 0.479 |
| BERTScore | 0.319 | 0.409 | 0.292 | 0.398 | 0.354 | 0.517 |
| (sent) | 0.310 | 0.517 | 0.427 | 0.517 | 0.387 | 0.537 |
| (doc) | 0.476 | 0.531 | 0.262 | 0.344 | 0.375 | 0.476 |
| (sent) | 0.329 | 0.371 | 0.341 | 0.460 | 0.345 | 0.459 |
| (doc) | 0.408 | 0.300 | 0.269 | 0.264 | 0.352 | 0.311 |
| FactScore (GPT4-o) | 0.365 | 0.511 | 0.362 | 0.460 | 0.405 | 0.549 |
| ARCscore (GPT-4o) | 0.499 | 0.607 | 0.401 | 0.482 | 0.465 | 0.593 |
| ARCscore (DeepSeek-R1) | 0.545 | 0.653 | 0.441 | 0.519 | 0.509 | 0.638 |
Appendix G Zero-shot summaries with argument-aware prompts
We perform a zero-shot ablation using four representative models from our set of eight: Llama-3.2-1B, Llama-3.2-3B, Qwen-2.5-3B, and Qwen-2.5-7B, varying both model families and sizes. In this setting, models are instructed to generate summaries of the same length as the human references, with an explicit emphasis on the key arguments in the input text. This experiment evaluates whether explicitly encoding argument salience in the prompt improves the inclusion of salient arguments in model-generated summaries. The zeroshot prompt used is described in Table 16.
| Argument-Aware Zero-Shot Prompt |
| Input: |
| Read the following text carefully: |
| {input} |
| Task Description: |
| Write an abstractive summary in about {length} words. |
| Instructions: |
| • Focus primarily on the key arguments presented in the text. • Ensure the summary is coherent, fluent, and written in continuous prose (not bullet points or key phrases). |
| Output: |
| Summary: |
Given the consistency across judges and datasets we report results on the smaller CANLII set (CANLII100) and the full DRI set using GPT-4-o as a judge.
Table 17 shows that explicitly instructing models to focus on key arguments did not consistently improve performance and, in some cases, even degraded coverage of salient roles. Only Qwen-2.5-3B,7B achieved a modest improvement on CANLII. These findings suggest that simply prompting models to emphasize arguments is insufficient for enhancing saliency coverage, underscoring the need for deeper alignment or fine-tuning strategies that more effectively encode argumentative relevance.
| Model | CANLII | DRI | ||
| Vanilla | Arg-Prompt | Vanilla | Arg-Prompt | |
| LLaMA-3.2-1B | 0.451 | 0.426 | 0.674 | 0.645 |
| LLaMA-3.2-3B | 0.573 | 0.570 | 0.705 | 0.667 |
| Qwen-2.5-3B | 0.648 | 0.654 | 0.730 | 0.718 |
| Qwen-2.5-7B | 0.676 | 0.703 | 0.799 | 0.754 |
Appendix H Source sentences identification
Algorithm 2 presents the lexical greedy procedure adopted from Ravaut et al. (2024) for identifying source sentences given a generated summary. We extend the original algorithm to return both the selected sentence indices and their corresponding argument role types, based on previously annotated sentence-level roles.
Appendix I Bias analysis for argument coverage without argument role normalization
While normalization in computing corrects for frequency skew, it may also understate coverage for dominant roles with inherently high raw scores. To provide a fuller picture, we additionally compute role-specific bias directly as . As shown in Figure 7, results on CANLII100 confirm prior findings: LLMs disproportionately prioritize conclusions over issues and reasons, both under length-controlled and non-controlled settings. For DRI, removing frequency normalization explains partially the higher scores reported in Section 5.4.
Appendix J Precision–recall trade-offs in ARC
ARC is designed to prioritize recall, aligning with prior coverage-based factuality metrics (e.g., FactScore) that emphasize capturing as many salient facts as possible in the summary. Nevertheless, we believe that incorporating precision-based analyses provides a more complete view of different models behavior. To this end, we conducted an additional study on summaries from the CANLII dataset, evenly sampled across summary length buckets. Using the same atomic fact decomposition and entailment filtering setup as in the main experiments, we compute: Precision: proportion of covered salient facts divided by the number of facts generated in the summary. Recall: proportion of reference salient facts covered by the summary. F1: harmonic mean of precision and recall. Both atomic fact decomposition and evaluation were conducted using GPT4-o.
Table 18 reports precision, recall, and F1 scores across models.
| Model | Precision | Recall | F1 |
| Qwen-2.5-14B | 0.8766 | 0.6418 | 0.7149 |
| Qwen-2.5-7B | 0.6337 | 0.6374 | 0.5737 |
| Qwen-2.5-3B | 0.5027 | 0.5874 | 0.5305 |
| Qwen-2.5-1.5B | 0.7733 | 0.5608 | 0.6227 |
| Mistral-8B | 0.7695 | 0.5188 | 0.5724 |
| LLaMA-3.1-8B | 0.7734 | 0.6174 | 0.6473 |
| LLaMA-3.2-3B | 0.5622 | 0.3983 | 0.4354 |
| LLaMA-3.2-1B | 0.4618 | 0.4069 | 0.4080 |
To further understand how evaluation choice affects model comparison, Table 19 shows model rankings induced by each metric.
| Model | Precision Rank | Recall Rank | F1 Rank |
| Qwen-2.5-14B | 1 | 1 | 1 |
| Qwen-2.5-7B | 5 | 3 | 5 |
| Qwen-2.5-3B | 6 | 5 | 4 |
| Qwen-2.5-1.5B | 3 | 4 | 3 |
| Mistral-8B | 4 | 6 | 6 |
| LLaMA-3.1-8B | 2 | 2 | 2 |
| LLaMA-3.2-3B | 7 | 8 | 7 |
| LLaMA-3.2-1B | 8 | 7 | 8 |
Finally, Table 20 reports rank correlation coefficients between rankings induced by different metrics.
| Metric Pair | Kendall | Pearson |
| Precision vs. Recall | 0.50 | 0.62 |
| Precision vs. F1 | 0.86 | 0.95 |
| Recall vs. F1 | 0.64 | 0.81 |
Overall, while precision introduces moderate reordering among models, the overall relative ranking of models remains largely stable. The positive correlations between recall and precision indicate that ARC’s recall-oriented design captures a general performance trend that largely persists even under precision-aware evaluation.